Two numbers are given. The sum of their squares minus twice their product is equal to 16. What is the difference between the largest and the smallest of these numbers?
(A) 4 (B) 6 (C) 8 (D) 10 (E) 16

Answer by mananth(16949)

Two numbers are given. The sum of their squares minus twice their product is equal to 16. What is the difference between the largest and the smallest of these numbers?

Let the two numbers be x & y
x^2+y^2-2xy = 16
(x-y)^2= 16
x-y =4
OR
-(x-y)=4
The difference =4